OpenStudy (anonymous):

Complete the following exercises by applying polynomial identities to complex numbers. Factor x2 + 64. Check your work. Factor 16x2 + 49. Check your work. Find the product of (x + 9i)2. Find the product of (x – 2i)2. Find the product of (x + (3+5i))2. PLEASE SOME HELP

OpenStudy (anonymous):

@alec.mccarty

OpenStudy (anonymous):

\[x^2+64\] has zeros that are easy to find \[x^2+64=0\\ x^2=-64\\ x=\pm\sqrt{-64}=\pm 8i\]

OpenStudy (anonymous):

therefore, over the complex numbers is factors as \[x^2+64=(x+8i)(x-8i)\]

OpenStudy (anonymous):

i love you thank you !